OpenStudy (anonymous):

Solve: 5(x + 6) 3x - 4 x -17 x 17 x -17 x -5

Parth (parthkohli):

The question does not make sense.

OpenStudy (anonymous):

5(x + 6) <= 3x - 4

OpenStudy (anonymous):

wait hold on let me rewrite it

Parth (parthkohli):

Hint: Distributive property to start with, then just solving like an equation.

Parth (parthkohli):

Pretend that the question is\[5(x + 6) = 3x - 4 \]How would you solve that?

OpenStudy (anonymous):

expand brackets then make x the subject

OpenStudy (anonymous):

Solve: 5(x + 6) <= 3x - 4 x >= -17 x >= 17 x <= -17 x <= -5

OpenStudy (anonymous):

start with 5x + 30 ≤ 3x - 4 and make x the subject

Parth (parthkohli):

Hmm, first tell what you get when you solve \(5(x + 6) = 3x - 4\)

OpenStudy (anonymous):

you have to multiple 5 ^ x and 5 ^ 6

Parth (parthkohli):

Yes, but what's "^" doing there?

OpenStudy (anonymous):

times

OpenStudy (anonymous):

I use it for times

Parth (parthkohli):

Okay, I get it. You are right... and then add those.

OpenStudy (anonymous):

5x + 30 = 3x - 4 subtract 3 to both sides -3x -3x 2x + 30 = -4 subtract 30 to both sides -30 -30 2x = -34 divide 2 to both sides 2 2 x = -17

OpenStudy (anonymous):

now change x = -17 to x <= -17 then flip the sign x >= -17

OpenStudy (anonymous):

correct????

Parth (parthkohli):

Why flip the sign? lol

OpenStudy (anonymous):

cause the multiplying

Parth (parthkohli):

You didn't multiply by any number.

OpenStudy (anonymous):

or not multiplying :PPP

OpenStudy (anonymous):

idk i forget

Parth (parthkohli):

No need to flip the sign; you got that right.

OpenStudy (anonymous):

i just thought u were supposed to flip the sign when you finish the equation

OpenStudy (anonymous):

ok thank you

Parth (parthkohli):

Only when you multiply by a negative.

OpenStudy (anonymous):

ok